OpenStudy (anonymous):

Completely simplify the following expressions. 1.

OpenStudy (anonymous):

y^9 z^2 xy^8 x ^-2 z^4

OpenStudy (anonymous):

the snitch

OpenStudy (anonymous):

what?

OpenStudy (mathstudent55):

When multiplying powers with the same base, add the exponents. Example: a^7 * a^-2 = a^5 Example: a^3 * b^5 * a^-3 * b^6 = a^0 * b^11 = b^11 (since a^0 = 1)

OpenStudy (anonymous):

???